# Can Diffusion Models Provide Rigorous Uncertainty Quantification for Bayesian Inverse Problems?

### v1

@@ 3.3 EXAMPLE ALGORITHMS
The BIPSDA framework also provides considerable flexibility regarding the choice of sampling scheme. As in DAPS [16], Markov chain Monte Carlo (MCMC) algorithms, such as Langevin dynamics or Hamiltonian Monte Carlo [36], can be employed. The prediction distribution can also be “sampled” by solving for the MAP point of the distribution, as in the DiffPIR algorithm [17]. Finally, in this work we also consider the use of a novel approach, inspired by the randomize-then-optimize (RTO) [18, 37, 19] algorithm, for approximate sampling from the prediction distribution. Here the key idea is to obtain an approximate sample from the prediction distribution by first adding noise to both the measurement and the denoising distribution, then solving for the MAP point of the corresponding noise-perturbed prediction distribution. Concretely, assuming a Gaussian approximation of the denoising distribution, noise is added to both the mean $\mathbf{m}_{\mathrm{aprx}}$ of this distribution and to the measurement $\mathbf{y}$ in the likelihood function. The noise-perturbed mean and measurement, denoted $\mathbf{m}_{\mathrm{aprx}}^{\prime}$ and $\mathbf{y}^{\prime}$ respectively, can therefore be written as follows:
@@
$\mathbf{m}_{\mathrm{aprx}}^{\prime}$ $=\mathbf{m}_{\mathrm{aprx}}+\mathbf{r}_{1},$
$\mathbf{y}^{\prime}$ $=\mathbf{y}+\mathbf{r}_{2}.$ (10)
@@
Here $\mathbf{r}_{1}\sim\mathcal{N}(\mathbf{r}_{1};\mathbf{0},\mathbf{C}_{\mathrm{aprx}})$, while $\mathbf{r}_{2}$ is sampled from the measurement noise distribution, e.g., for the additive Gaussian noise model in (1), we have $\mathbf{r}_{2}\sim\mathcal{N}(\mathbf{r}_{2};\mathbf{0},\boldsymbol{\Sigma}_{\mathbf{z}})$. This approach, like the MAP approach used in DiffPIR, has the advantage of enabling efficient MAP solvers to be employed for sampling from prediction distribution. Further, in the linear-Gaussian likelihood setting, where the corresponding prediction distribution is also Gaussian, it can be proven that this approach corresponds to exact sampling from the prediction distribution [18, 38].
@@ 4 NUMERICAL STUDIES
The performance of diffusion model based posterior sampling algorithms is often evaluated on high-dimensional problems (e.g., imaging problems), where the prior distribution is characterized implicitly through a dataset of examples (samples). In this setting, the posterior density is analytically unknown, making it difficult to evaluate the ability of the algorithm under consideration to capture the global structure of the posterior distribution and facilitate uncertainty quantification. To address this issue, we created a suite of benchmark problems for which the prior distribution, as well as the score of the denoising distribution, is analytically known. Using these benchmark problems, which use a Gaussian mixture as the prior distribution, we can rigorously assess the performance of the algorithms in our framework, which include the popular DAPS [16] and DiffPIR [17] algorithms.
The use of a Gaussian mixture prior to test the performance of diffusion model based posterior samplers has a number of advantages. First, in this setting the posterior density function is known, and in certain cases (e.g., linear-Gaussian likelihood function, where the posterior is also a Gaussian mixture) can be straightforwardly sampled from, enabling us to investigate errors in the global structure of the samples from these algorithms. Second, under this choice $\pi_{t}(\mathbf{m}(t))$ is also a Gaussian mixture and the noisy prior score $\nabla_{\mathbf{m}(t)}\log\pi_{t}(\mathbf{m}(t))$ is known analytically. This enables the decoupling of error in the score modeling error from error inherent to the sampling algorithm under consideration in a principled manner. Third, Gaussian mixture distributions can be made arbitrarily complex through the choice of the number of modes and the covariances of each component Gaussian.
@@
In each study, we tested the performance of nine algorithms within the BIPSDA framework, with each algorithm characterized by the method used to approximate the denoising distribution and the algorithm used to sample from the prediction distribution. Additional details regarding the implementation of these algorithms, as well as the problem setting, and metrics used to evaluate the performance of the algorithms, are provided in the following subsections.
@@ 4.1 PROBLEM SETTINGS
where each $\pi_{\mathrm{pr},i}$ is a Gaussian, $w_{i}$ is the weight of the $i$th component, and we set $N_{m}=3$. Here $\pi_{\mathrm{pr},1}$ has mean $[-5,\cdots,-5]^{T}$ and identity covariance matrix, $\pi_{\mathrm{pr},2}$ has mean $[0,\cdots,0]^{T}$ and a diagonal covariance matrix with entries linearly spaced between $1$ and $2$, and $\pi_{\mathrm{pr},3}$ has mean $[5,\cdots,5]^{T}$ and a covariance matrix with the same eigenvalues as $\pi_{\mathrm{pr},2}$ but randomly chosen eigenvectors. The corresponding weights were set as $w_{1}=0.4$, $w_{2}=0.3$, and $w_{3}=0.3$. A plot of two components of samples from this distribution is shown in Figure 1.
@@
where $\mathrm{Poi}(\cdot)$ is the Poisson distribution, the exponential operator is applied pointwise, and $\mathbf{C}\in\mathbb{R}^{15\times 10}$. We set $I_{0}=1000$ and populated $\mathbf{C}$ with i.i.d. random entries from the uniform distribution over the interval [.01,.05].01.05[.01,.05][ .01 , .05 ] ($\mathrm{SNR}\approx 23.1709\;\mathrm{dB}$).
@@ 4.2 ALGORITHM IMPLEMENTATIONS
As previously discussed, in each study we analyzed the performance of nine BIPSDA algorithms, with each algorithm corresponding to a different combination of the ODE, TU, and TC methods for approximating the denoising distribution, and the Langevin dynamics (‘Lang’), MAP estimation (‘MAP’), or RTO based approaches for sampling from the corresponding prediction distribution. Throughout the remainder of this paper, we denote each of the nine BIPDSA algorithms by the name of the sampling algorithm and the name of the method for approximating the denoising distribution. For example, the BIPDA algorithm that uses Langevin dynamics to sample from the prediction distribution and the TU method to approximate the denoising distribution is denoted Lang-TU.
Each of the BIPSDA algorithms tested was implemented using $N_{A}=200$ timepoints in the noise annealing, with the $t_{i}$ set by polynomial interpolation between $T$ and 00. The ODE method for approximating the denoising distribution was implemented by solving the probability flow ODE using the Euler discretization with five discretization time-steps. The above settings are the same as used by Zhang et al; see [16] for details.
The implementation of the Langevin dynamics based sampler and the MAP solver used in the MAP and RTO-based approaches was problem specific. The sampler for the stylized inpainting problem was implemented with step size $=5\times 10^{-5}$, $100$ subiterations per algorithm iteration, and no Metropolis correction, as in [16]. The implementation was the same for the stylized x-ray tomography based problem, but with $1000$ subiterations instead of $100$. For the phase retrieval problem, we observed severe stability issues with unadjusted Langevin dynamics, and instead implemented Langevin dynamics with Metropolis adjustment and preconditioning [41]. Here $1000$ subiterations were used with step size =.$=.2$, and at each iteration we fixed the preconditioning matrix as the average of a set of matrices, where each matrix in the set is the Gauss-Newton Hessian approximation [42] evaluated at different representative sample points. Regarding the MAP solver used in the MAP and RTO-based approaches, for the stylized inpainting problem we exploited the fact that the MAP point has a closed-form expression to compute the MAP point exactly. For the stylized x-ray tomography and phase retrieval problems, we estimated the MAP point using the PyTorch implementation of the limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BFGS) algorithm [42] (40 subiterations per algorithm iteration, strong Wolfe line search).
[…]
@@ 4.3 EVALUATION METHODS
$=\sum_{i=1}^{N_{m}}w_{i}\,\pi_{\mathrm{like}}(\mathbf{y}\mid\mathbf{m})\,\pi_{\mathrm{pr},i}(\mathbf{m})$
@@
$\mathrm{CMD}(\pi_{1},\pi_{2})$ $=\frac{1}{\alpha}\left\lVert\mathbb{E}_{\pi_{1}}[\mathbf{m}]-\mathbb{E}_{\pi_{2}}[\mathbf{m}]\right\rVert_{2}$
@@
$\eta_{\mathrm{max}}=\mathbb{E}_{\mathbf{y}}\left[||\boldsymbol{\eta}(\mathbf{y})||_{\infty}\right]$
@@
The MMD metric is based on the difference between the two probability distributions in a reproducing kernel Hilbert space (RKHS) [21]. Here we chose the kernel $k(\cdot,\cdot)$ in the metric to be the sum of Gaussian radial basis functions with different bandwidths, i.e.,
$k(\mathbf{m}_{1},\mathbf{m}_{2})=\sum_{i=1}^{N_{b}}\mathrm{exp}\left(-\frac{||\mathbf{m}_{1}-\mathbf{m}_{2}||_{2}^{2}}{\epsilon_{i}}\right),$
@@
Mean Variance CMD MMD Mean Variance CMD MMD
Lang-TU $0.029$ $(0.011,0.055)$ $0.64$ $(0.44,0.86)$ $0.04$ $(0.03,0.06)$ $0.043$ $(0.033,0.054)$ $0.046$ $(0.013,0.092)$ $0.64$ $(0.43,0.87)$ $0.05$ $(0.03,0.07)$ $0.043$ $(0.032,0.055)$
Lang-TC $0.024$ $(0.009,0.048)$ $0.64$ $(0.43,0.87)$ $0.04$ $(0.03,0.06)$ $0.043$ $(0.033,0.052)$ – – – – – – – –
Lang-ODE $0.032$ $(0.016,0.060)$ $0.10$ $(0.02,0.21)$ $\mathbf{0.01}$ $(\mathbf{0.01},\mathbf{0.02})$ $0.002$ $(0.001,0.003)$ $0.047$ $(0.016,0.090)$ $0.11$ $(0.02,0.25)$ $0.02$ $(0.01,0.03)$ $0.002$ $(0.001,0.005)$
MAP-TU $\mathbf{0.018}$ $(\mathbf{0.007},\mathbf{0.031})$ $0.95$ $(0.71,1.19)$ $0.06$ $(0.04,0.08)$ $0.142$ $(0.135,0.149)$ $0.046$ $(0.013,0.093)$ $0.95$ $(0.70,1.21)$ $0.07$ $(0.04,0.09)$ $0.143$ $(0.133,0.155)$
MAP-TC $\mathbf{0.018}$ $(\mathbf{0.007},\mathbf{0.030})$ $0.96$ $(0.71,1.20)$ $0.06$ $(0.04,0.08)$ $0.143$ $(0.135,0.151)$ – – – – – – – –
MAP-ODE $0.023$ $(0.013,0.037)$ $0.47$ $(0.34,0.61)$ $0.03$ $(0.02,0.04)$ $0.028$ $(0.025,0.032)$ $0.047$ $(0.016,0.090)$ $0.47$ $(0.33,0.64)$ $0.04$ $(0.02,0.06)$ $0.029$ $(0.025,0.035)$
RTO-TU $0.022$ $(0.008,0.037)$ $0.05$ $(0.02,0.07)$ $\mathbf{0.01}$ $(\mathbf{0.01},\mathbf{0.01})$ $\mathbf{0.001}$ $(\mathbf{0.000},\mathbf{0.001})$ $\mathbf{0.044}$ $(\mathbf{0.011},\mathbf{0.089})$ $\mathbf{0.06}$ $(\mathbf{0.02},\mathbf{0.09})$ $\mathbf{0.01}$ $(\mathbf{0.01},\mathbf{0.02})$ $\mathbf{0.001}$ $(\mathbf{0.001},\mathbf{0.003})$
RTO-TC $0.020$ $(0.008,0.031)$ $\mathbf{0.03}$ $(\mathbf{0.01},\mathbf{0.06})$ $\mathbf{0.01}$ $(\mathbf{0.00},\mathbf{0.01})$ $\mathbf{0.001}$ $(\mathbf{0.000},\mathbf{0.001})$ – – – – – – – –
RTO-ODE $0.027$ $(0.013,0.044)$ $1.05$ $(0.79,1.28)$ $0.08$ $(0.05,0.10)$ $0.051$ $(0.047,0.056)$ $0.046$ $(0.016,0.090)$ $1.01$ $(0.80,1.22)$ $0.08$ $(0.06,0.10)$ $0.050$ $(0.043,0.057)$
Reference $0.020$ $(0.009,0.034)$ $0.03$ $(0.01,0.07)$ $0.01$ $(0.00,0.01)$ $0.001$ $(0.000,0.001)$ $0.020$ $(0.009,0.034)$ $0.03$ $(0.01,0.07)$ $0.01$ $(0.00,0.01)$ $0.001$ $(0.000,0.001)$
Table 1: Stylized inpainting study, low noise regime: Average errors in estimation of the mean and pointwise variance of the posterior distribution, and average central moment discrepancy (CMD) and maximum mean discrepancy (MMD) errors, for different BIPSDA methods (with both learned and analytic score). Interdecile ranges are also displayed in parentheses, and a reference that shows the average errors between two sets of i.i.d. samples from the ground truth posterior is included. The best-performing method with respect to each error measure is bolded. In the learned score case, with the exception of the Tweedie-Correlated (TC) approaches that were not implemented, all approaches performed well, with the MAP and RTO based approaches providing particularly strong performance.

### v2

@@ 3.3.2 Sampling from the prediction distribution
Finally, the ‘RTO’ variant is a novel contribution of our work. It is inspired by the randomize-then-optimize (RTO) [24, 26, 25] algorithm for approximate sampling from the prediction distribution. Here the key idea is to obtain an approximate sample from the prediction distribution by first adding noise to both the measurement and the denoising distribution mean, then solving for the MAP point of the corresponding noise-perturbed prediction distribution. Concretely, assuming the Gaussian approximation of the denoising distribution in Eq. (9) and the Gaussian additive noise model in Eq. (1), ‘RTO’ generates an approximate sample $\mathbf{m}(0)$ from the prediction distribution by solving
@@
$\mathbf{m}(0)=\operatornamewithlimits{argmax}_{\mathbf{m}}\pi_{\rm like}(\mathbf{y}^{\prime}|\mathbf{m})\;\mathcal{N}(\mathbf{m};\mathbf{m}_{\mathrm{aprx}}^{\prime},\mathbf{C}_{\mathrm{aprx}}),$ (10)
where
$\mathbf{m}_{\mathrm{aprx}}^{\prime}$ $\sim\mathcal{N}(\mathbf{m}_{\mathrm{aprx}}^{\prime};\mathbf{m}_{\mathrm{aprx}},\mathbf{C}_{\mathrm{aprx}}),\text{ and}$
$\mathbf{y}^{\prime}$ $\sim\mathcal{N}(\mathbf{y}^{\prime};\mathbf{y},\boldsymbol{\Sigma}_{\mathbf{z}})$
@@
denote the noise-perturbed mean and measurement, respectively. Note that this approach can be viewed as an application of Algorithm 1 in [25] to the subproblem of sampling from the prediction distribution (line 6 in Algorithm 1). Like the ‘MAP’ variant used in DiffPIR, it has the advantage of enabling efficient MAP solvers to be employed for sampling from prediction distribution. Further, in the linear-Gaussian likelihood setting, where the corresponding prediction distribution is also Gaussian, it can be proven that this approach corresponds to exact sampling from the prediction distribution [24, 43].
3.3.3 Analyzed variants
Mixing and matching different variants to approximate the denoising distribution with those to sample from the prediction distribution, we obtain the 9 variants summarized in Table 1. The DAPS and DiffPIR algorithms correspond to the ‘Lang-ODE’ and ‘MAP-TU’ variants of our framework, respectively.
Denoising Dist. Approximation
   ODE    TU    TC
Sampling Alg. Lang Lang-ODE Lang-TU Lang-TC
MAP MAP-ODE MAP-TU MAP-TC
RTO RTO-ODE RTO-TU RTO-TC
Table 1: Visual representation of the nine example algorithms tested, each of which corresponds to different choices of the denoising distribution approximation (line 5 in 1) and sampling scheme (line 6 in 1). Note that ‘Lang-ODE’ is the DAPS algorithm, while ‘MAP-TU’ is the DiffPIR algorithm; the other seven variants are novel.
@@ 4 NUMERICAL STUDIES
The performance of diffusion-model-based posterior sampling algorithms is often evaluated on high-dimensional problems (e.g., imaging problems), where the prior distribution is characterized implicitly through a dataset of examples (samples). In this setting, the posterior density is analytically unknown, making it difficult to evaluate the ability of the algorithm under consideration to capture the global structure of the posterior distribution and facilitate uncertainty quantification. To address this issue, we created a suite of benchmark problems for which the prior distribution, as well as the score of the denoising distribution, is analytically known. Using these benchmark problems, which use a Gaussian mixture as the prior distribution, we can rigorously assess the performance of the algorithms in our framework, which include the popular DAPS [22] and DiffPIR [23] algorithms.
The use of a Gaussian mixture prior to test the performance of diffusion-model-based posterior samplers has a number of advantages. First, in this setting the posterior density function is known, and in certain cases (e.g., linear-Gaussian likelihood function, where the posterior is also a Gaussian mixture) can be straightforwardly sampled from, enabling us to investigate errors in the global structure of the samples from these algorithms. Second, under this choice $\pi_{t}(\mathbf{m}(t))$ is also a Gaussian mixture and the noisy prior score $\nabla_{\mathbf{m}(t)}\log\pi_{t}(\mathbf{m}(t))$ is known analytically. This enables the decoupling of error in the score modeling error from error inherent to the sampling algorithm under consideration in a principled manner. Third, Gaussian mixture distributions can be made arbitrarily complex through the choice of the number of modes and the covariances of each component Gaussian.
@@
In each study, we tested the performance of the nine algorithms within the BIPSDA framework summarized in Table 1. Additional details regarding the implementation of these algorithms, as well as the problem setting and metrics used to evaluate the performance of the algorithms, are provided in the following subsections.
@@ 4.1 PROBLEM SETTINGS
where each $\pi_{\mathrm{pr},i}$ is a Gaussian, $w_{i}$ is the weight of the $i$th component, and we set $N_{m}=3$. Here $\pi_{\mathrm{pr},1}$ has mean $5,\cdots,-5]^{T}$ and identity covariance matrix, $\pi_{\mathrm{pr},2}$ has mean $0,\cdots,0]^{T}$ and a diagonal covariance matrix with entries linearly spaced between $1$ and $2$, and $\pi_{\mathrm{pr},3}$ has mean $5,\cdots,5]^{T}$ and a covariance matrix with the same eigenvalues as $\pi_{\mathrm{pr},2}$ but randomly chosen eigenvectors. The corresponding weights were set as $w_{1}=0.4$, $w_{2}=0.3$, and $w_{3}=0.3$. A plot of two components of samples from this distribution is shown in Figure 1.
@@
where $\mathrm{Poi}(\cdot)$ is the Poisson distribution, the exponential operator is applied pointwise, and $\mathbf{C}\in\mathbb{R}^{15\times 10}$. We set $I_{0}=1000$ and populated $\mathbf{C}$ with i.i.d. random entries from the uniform distribution over the interval $[.01,.05]$ ($\mathrm{SNR}\approx 23.1709\;\mathrm{dB}$).
@@ 4.2 ALGORITHM IMPLEMENTATIONS
Each of the BIPSDA algorithms variants summarized in Table 1 was implemented using $N_{A}=200$ timepoints in the noise annealing, with the $t_{i}$ set by polynomial interpolation between $T$ and 0. The ‘ODE’ and ‘TU’ variants for approximating the denoising distribution were implemented with $\beta(t)=\sigma(t)$ in the covariance approximation $\mathbf{C}_{\mathrm{aprx}}=\beta(t)^{2}\mathbf{I}$; the ‘ODE’ method for approximating the denoising distribution was implemented by solving the probability flow ODE using the Euler discretization with five discretization time-steps. The above settings are the same as used by Zhang et al; see [22] for details.
The implementation of the Langevin dynamics based sampler and the MAP solver used in the ‘MAP‘ and ‘RTO‘ variants was problem specific. The sampler for the stylized inpainting problem was implemented with step size $5\times 10^{-5}$, $100$ subiterations per algorithm iteration, and no Metropolis correction, as in [22]. The implementation was the same for the stylized x-ray tomography based problem, but with $1000$ subiterations instead of $100$. For the phase retrieval problem, we observed severe stability issues with unadjusted Langevin dynamics, and instead implemented Langevin dynamics with Metropolis adjustment and preconditioning [46]. Here $1000$ subiterations were used with step size =.2=.2, and at each iteration we fixed the preconditioning matrix as the average of a set of matrices, where each matrix in the set is the Gauss-Newton Hessian approximation [47] evaluated at different representative sample points. Regarding the MAP solver used in the ‘MAP‘ and ‘RTO‘ variants, for the stylized inpainting problem we exploited the fact that the MAP point has a closed-form expression. For the stylized x-ray tomography and phase retrieval problems, we estimated the MAP point using the PyTorch implementation of the limited-memory Broyden–Fletcher–Goldfarb–Shanno (L-BFGS) algorithm [47] (40 subiterations per algorithm iteration, strong Wolfe line search).
Each of the algorithms described above was implemented using the same pretrained score model corresponding to a diffusion model with $\sigma(t)=t$ and $T=10$. The score model was given as $s_{\boldsymbol{\theta}}(\mathbf{m}(t),t)=\nabla_{\mathbf{m}(t)}g_{\boldsymbol{\theta}}(\mathbf{m}(t),t)$ where $g_{\boldsymbol{\theta}}(\mathbf{m}(t),t):\mathbb{R}^{10}\times\mathbb{R}_{+}\to\mathbb{R}$ was a deep neural network. In particular, the network architecture had six total hidden layers (width =512=512), with the noise level $\sigma(t)$ appended to the input $\mathbf{m}(t)$ and to the network state after the 4th layer, width =512=512. The network weights were trained to minimize the loss function in (6), with $N=80,000$ total samples in the training set, $w(t)=\sigma^{2}(t)$, and $T_{\mathrm{min}}=.01$. The Adam optimizer [48] was used with learning rate $10^{-5}$ and a batch size of $8000$ $\{\mathbf{m}_{i},\mathbf{z}_{i},t_{i}\}$ pairs (with the $\mathbf{z}_{i}$ and $t_{i}$ picked i.i.d. at each iteration). The network was trained for $50,000$ optimization iterations. Additionally, each algorithm was also tested with the exact ground truth score $\nabla_{\mathbf{m}(t)}\log\pi_{t}\mathbf{m}(t)$ used instead of the score model to isolate the effect of the score modeling error from other sources of error in the algorithms’ performance.
@@ 4.3 EVALUATION METHODS
$=\sum_{i=1}^{N_{m}}w_{i}\,\pi_{\mathrm{like}}(\mathbf{y}\mid\mathbf{m})\,\pi_{\mathrm{pr},i}(\mathbf{m}),$ (11)
@@
We assessed the convergence of the MCMC samples to the target distribution using well-established MCMC diagnostics, including the potential scale reduction factor (PSRF) [52] and effective sample size (ESS) [53] metrics. These metrics were computed for a randomly selected single posterior sampling trial. For the x-ray tomography problem, the maximum PSRF value over all parameter dimensions was 1.0006, with PSRF values less than 1.01 generally indicating that the parallel chains have stabilized, and samples are likely to have reached the target distribution [52]. The minimum ESS computed for each parameter independently was $47,077.65$, which is much larger than the number of posterior samples used in our evaluation metrics ($10,000$). For the phase retrieval study, PSRF and ESS were computed separately for each PyDREAM run corresponding to each component of the posterior mixture in Eq. (4.3). The maximum PSRF value was $1.0004$ and the minimum ESS was $49,862.49$, which again indicate that a sufficient number of high-quality samples from the target distribution were generated.
@@
$\mathrm{CMD}(\pi_{1},\pi_{2})$ $=\frac{1}{\alpha}\left\lVert\mathbb{E}_{\pi_{1}}[\mathbf{m}]-\mathbb{E}_{\pi_{2}}[\mathbf{m}]\right\rVert_{2}$ (12)
@@
[…]
@@
The MMD metric is based on the difference between the two probability distributions in a reproducing kernel Hilbert space (RKHS) [28]. In particular, the MMD metric is an integral probability metric in the RKHS space and can be written as
$\mathrm{MMD}(\pi_{1},\pi_{2})=\sup_{f\in\mathcal{F}}\left(\mathbb{E}_{\pi_{1}}[f(\mathbf{m})]-\mathbb{E}_{\pi_{2}}[f(\mathbf{m})]\right),$ (13)
where $\pi_{1}$ and $\pi_{2}$ are probability distributions and $\mathcal{F}$ is the set of functions lying on the unit ball in the RKHS. An RKHS space is characterized by the choice of kernel function $k(\cdot,\cdot)$, with $\{k(\cdot,\mathbf{m})\mid\mathbf{m}\in\mathbb{R}^{D}\}$ forming a set of basis functions for $\mathcal{F}$. In this work, we set the kernel as the sum of Gaussian radial basis functions with different bandwidths, i.e.,
@@
with $N_{b}=5$, $\epsilon_{i}=\bar{\epsilon}\;2^{i-\lceil N_{b}/2\rceil}$, and $\bar{\epsilon}$ set as the average squared two-norm distance between samples from the reference distribution (as in [28]).
@@
Mean Error Variance Error CMD MMD Mean Error Variance Error CMD MMD
Lang-TU
$0.029$
$(0.011,0.055)$
$0.64$
$(0.44,0.86)$
$0.04$
$(0.03,0.06)$
$0.043$
$(0.033,0.054)$
$0.046$
$(0.013,0.092)$
$0.64$
$(0.43,0.87)$
$0.05$
$(0.03,0.07)$
$0.043$
$(0.032,0.055)$
Lang-TC
$0.024$
$(0.009,0.048)$
$0.64$
$(0.43,0.87)$
$0.04$
$(0.03,0.06)$
$0.043$
$(0.033,0.052)$
Lang-ODE
$0.032$
$(0.016,0.060)$
$0.10$
$(0.02,0.21)$
$\mathbf{0.01}$
$(\mathbf{0.01},\mathbf{0.02})$
$0.002$
$(0.001,0.003)$
$0.047$
$(0.016,0.090)$
$0.11$
$(0.02,0.25)$
$0.02$
$(0.01,0.03)$
$0.002$
$(0.001,0.005)$
MAP-TU
$\mathbf{0.018}$
$(\mathbf{0.007},\mathbf{0.031})$
$0.95$
$(0.71,1.19)$
$0.06$
$(0.04,0.08)$
$0.142$
$(0.135,0.149)$
$0.046$
$(0.013,0.093)$
$0.95$
$(0.70,1.21)$
$0.07$
$(0.04,0.09)$
$0.143$
$(0.133,0.155)$
MAP-TC
$\mathbf{0.018}$
$(\mathbf{0.007},\mathbf{0.030})$
$0.96$
$(0.71,1.20)$
$0.06$
$(0.04,0.08)$
$0.143$
$(0.135,0.151)$
MAP-ODE
$0.023$
$(0.013,0.037)$
$0.47$
$(0.34,0.61)$
$0.03$
$(0.02,0.04)$
$0.028$
$(0.025,0.032)$
$0.047$
$(0.016,0.090)$
$0.47$
$(0.33,0.64)$
$0.04$
$(0.02,0.06)$
$0.029$
$(0.025,0.035)$
RTO-TU
$0.022$
$(0.008,0.037)$
$0.05$
$(0.02,0.07)$
$\mathbf{0.01}$
$(\mathbf{0.01},\mathbf{0.01})$
$\mathbf{0.001}$
$(\mathbf{0.000},\mathbf{0.001})$
$\mathbf{0.044}$
$(\mathbf{0.011},\mathbf{0.089})$
$\mathbf{0.06}$
$(\mathbf{0.02},\mathbf{0.09})$
$\mathbf{0.01}$
$(\mathbf{0.01},\mathbf{0.02})$
$\mathbf{0.001}$
$(\mathbf{0.001},\mathbf{0.003})$
RTO-TC
$0.020$
$(0.008,0.031)$
$\mathbf{0.03}$
$(\mathbf{0.01},\mathbf{0.06})$
$\mathbf{0.01}$
$(\mathbf{0.00},\mathbf{0.01})$
$\mathbf{0.001}$
$(\mathbf{0.000},\mathbf{0.001})$
RTO-ODE
$0.027$
$(0.013,0.044)$
$1.05$
$(0.79,1.28)$
$0.08$
$(0.05,0.10)$
$0.051$
$(0.047,0.056)$
$0.046$
$(0.016,0.090)$
$1.01$
$(0.80,1.22)$
$0.08$
$(0.06,0.10)$
$0.050$
$(0.043,0.057)$
Reference
$0.020$
$(0.009,0.034)$
$0.03$
$(0.01,0.07)$
$0.01$
$(0.00,0.01)$
$0.001$
$(0.000,0.001)$
$0.020$
$(0.009,0.034)$
$0.03$
$(0.01,0.07)$
$0.01$
$(0.00,0.01)$
$0.001$
$(0.000,0.001)$
Table 2: Stylized inpainting study, low noise regime: Average errors in estimation of the mean and pointwise variance of the posterior distribution, and average central moment discrepancy (CMD) and maximum mean discrepancy (MMD) errors, for different BIPSDA methods (with both learned and analytic score). Interdecile ranges are also displayed in parentheses, and a reference that shows the average errors between two sets of i.i.d. samples from the ground truth posterior is included. The best-performing method with respect to each error measure is bolded. In the learned score case, with the exception of the Tweedie-Correlated (‘TC’) approaches that were not implemented, all approaches performed well, with the ‘MAP’ and ‘RTO’ variants providing particularly strong performance.